# Benchmarking Ultra-Low-Power $\mu$NPUs

### v1

@@ Abstract.
In this paper we present the first comparative evaluation of a number of commercially-available $\mu$NPUs, as well as the first independent benchmarks for several of these platforms. We develop and open-source a model compilation framework to enable consistent benchmarking of quantized models across diverse $\mu$NPU hardware. Our benchmark targets end-to-end performance and includes model inference latency, power consumption, and memory overhead, alongside other factors. The resulting analysis uncovers both expected performance trends as well as surprising disparities between hardware specifications and actual performance, including $\mu$NPUs exhibiting unexpected scaling behaviors with increasing model complexity. Our framework provides a foundation for further evaluation of $\mu$NPU platforms alongside valuable insights for both hardware designers and software developers in this rapidly evolving space.
@@ 3.1. Hardware
Figure 2. a visualization of the various $\mu$NPUs used in our benchmark, and how they compare in terms of GOPS, peak power draw, and efficiency (GOPS/mW).
@@ 3.2. Models
CIFAR10-NAS: the optimal CNN model generated by neural architecture search (NAS) for the CIFAR-10 dataset, combining diverse convolutional units, pooling layers, and unique connectivity patterns. The model was generated using the Once-for-All (OFA) NAS framework, a weight-sharing-based framework that decouples search and training by constructing a supernet model from which various hardware-specific subnet models can be derived (onceforall, ). This is our largest model, with 74.3 MMACs and 36.4 MFLOPs. Trained on the CIFAR-10 dataset, with 3x32x32 input size and 10-class output.
@@ 3.2.1. Ensuring Model Uniformity
We encountered substantial variability in operator support across the benchmark platforms. The NXP-MCXN947’s eIQ Neutron NPU lacks native support for softmax operations, for example, necessitating its implementation as a CPU post-processing step for relevant models. Similarly, operations associated with NMS in the YoloV1 model were inconsistently supported across platforms, requiring us to also move the entire NMS operation to CPU post-processing. This explains the unusual multi-component output shape of our YoloV1 model (see Table 2). The benchmark platforms also outline varying levels of support for operator compatibility. The MAX78000, for example, only supports 1D convolution with kernel sizes 1 to 9 and 2D convolution with kernel sizes of 1 by 1 or 3 by 3. Unsupported operations will fall back to CPU execution and incur latency penalties.
@@ 3.3. Evaluation Metrics
3.4. Performance Evaluation
and Stage Breakdown
@@ 3.4. Performance Evaluation
NAS 1.10±0.002 0.85±0.001 0.10±0.002 0.31±0.002 0.79±0.007 0.83±0.006 0.01±0.001 0.03±0.001 0.01±0.001
ResNet 1.24±0.003 0.85±0.002 0.17±0.002 0.52±0.003 0.85±0.006 0.84±0.019 0.01±0.001 0.06±0.001 0.02±0.001
SimpleNet 2.29±0.006 1.65±0.003 0.16±0.005 0.61±0.004 0.89±0.006 0.99±0.006 0.01±0.001 0.07±0.001 0.02±0.001
Autoencoder 3.92±0.014 2.75±0.008 1.12±0.028 0.45±0.002 3.57±0.035 3.06±0.038 0.01±0.001 3.48±0.082 0.32±0.001
YoloV1_small 2.27±0.004 1.76±0.003 0.20± 0.005 0.44±0.006 0.73±0.009 0.81±0.008 0.01±0.001 0.05±0.001 0.01±0.001
@@
CIFAR10-NAS 1.11±0.002 0.85±0.001 0.11±0.002 0.31±0.002 0.98±0.009 1.04±0.008 2.75±0.281
ResNet 1.24±0.003 0.85±0.002 0.22±0.001 0.52±0.002 1.08±0.008 1.05±0.024 9.31±2.567
SimpleNet 2.30±0.006 1.66±0.003 0.21±0.007 0.62±0.004 1.13±0.008 1.29±0.010 4.13±0.459
Autoencoder 3.94±0.014 2.78±0.008 6.25±0.203 13.74±1.956 22.45±0.392 12.97±0.232 0.01±0.001
YoloV1_small 2.27±0.004 1.76±0.003 0.23±0.005 0.44±0.007 0.91±0.013 1.03±0.011 5.75±0.770
@@ 4. RESULTS & DISCUSSION
Table Supplementary Material, which can found in the supplementary material, details our full latency (ms) and power (mW) measurements across each stage, model, and platform.
@@ 4.1. Power and Efficiency Breakdown
The MAX78000 (C) with Cortex-M4 CPU active demonstrates the best efficiency across all evaluated models when NPU initialization overhead is considered, with consistent ¡30ms end-to-end latency. This aligns with previous standalone benchmarks (moss2022ultra, ).
@@
The MAX78000 (R) with RISC-V CPU active is next best, followed by the power-hungry but low-latency HX-WE2 platform, with Arm Corstone-300 (Cortex-M55 & Ethos-U55 NPU). Notably, the HX-WE2 beats the MAX78000 (R) in terms of efficiency for certain models, due to the MAX78000’s large memory I/O overhead. The HX-WE2 (S/P) demonstrates average $\sim$1.93x and $\sim$3.07x speedup in end-to-end latency over the MAX78000 (C) and (R) respectively, but $\sim$3.13x and $\sim$3.33x increase in average power consumption. We find the Vela Performance-optimized models, for the HX-WE2, generally achieve slightly lower latency than the Size-optimized models. However, their efficiency gain diminishes with model complexity – efficiency on Performance-optimized YoloV1 is lower than on its Size-optimized variant.
@@
The general-purpose MCUs without dedicated neural hardware – the STM32H7A3ZI and ESP32s3 – demonstrate significantly lower efficiency across all models. This result empirically validates the advantage neural hardware provides for performing on-device inference in constrained environments, with up to 2 orders of magnitude improvement in end-to-end latency in some cases. However, the STM32H7A3ZI’s power consumption during inference (54.91 - 56.11 mW) is comparable to or lower than MAX78000 (C)/(R) for some models. This is particularly evident for the Autoencoder model, where the STM32H7A3ZI achieves a surprisingly competitive 3.483 $I_{mJ}$ – comparable to the best-performing platforms in our suite. This anomaly is likely attributable to the STM32H7A3ZI’s relatively efficient Cortex-M7 core when operating on the Autoencoder’s simple computational structure (0.5455 MMACs). In contrast, the ESP32 consistently exhibits high inference power consumption (129.74 - 157.17 mW) and latency (7.11 – 536.22 ms), despite its advertised support for CPU-accelerated tensor operations. Altogether, while general-purpose MCUs can achieve reasonable efficiency for simple models, they quickly become impractical for more complex NNs.
@@
The MILK-V, our RISC-V SoC, also demonstrates low efficiency across all models, due to its NPU initialization overhead. We observe a different story, however, if initialization overhead is removed from consideration ($i.e.$, for continuous operation). Without initialization, the MILK-V ranks highest for efficiency across almost all benchmark models, averaging a 3.58x $I_{mJ}$ increase over the next best platform for each model. Notably, despite a large idle power draw, it achieves blazingly fast inference times (0.17 - 0.61 ms). Thus, for applications where power consumption isn’t a major constraint and workloads don’t require frequent NPU initialization/deinitialization, but low-latency and a compact form factor are key, the MILK-V could prove effective.
@@
Fig. 7 details the power consumption breakdown across all evaluated platforms; among these, the NXP-MCXN947 exhibits the largest overall power draw (105.71 – 118.57 mW), without much variation across the benchmark models.
@@
Beyond peak power, idle power consumption is another key consideration for low-power deployments, particularly if workloads run infrequently – idle power also varies significantly across our benchmark platforms. The MAX78000 demonstrates the lowest idle power of the various $\mu$NPU platforms (10.87 mW with RISC-V and 13.21 mW under Cortex-M4), with the GAP8 next (33.67 mW). The HX-WE2 and NXP-MCXN947 platforms rank highest (89.09 mW and 105.71 mW, respectively), raising concerns about their applicability in extremely power-constrained scenarios (such as ones in which long idle durations dominate overall energy usage).
@@ 4.2.2. Memory I/O
The significant memory I/O latency across all models on the MAX78000 forms an obvious inference bottleneck, with an average of 6.10x and 9.80x (Cortex-M4 and RISC-V) longer spent on memory I/O than actual inference ($e.g.$, 44.89 ms vs. 2.96 ms for ResidualNet with the MAX78k (R), meaning over 90% of end-to-end inference time is dedicated to memory operations rather than computation). This implies the MAX78000’s performance is largely memory-bound, and aligns with previous standalone benchmarks (moss2022ultra, ). Notably, memory I/O operations are more efficient on the MAX78000’s Cortex-M4 CPU than its RISC-V one.
@@
Memory I/O operations introduce negligible overhead across the other benchmark platforms with shared SRAM. The NXP-MCXN947 in particular features extremely low memory I/O overhead – as low as 0.05 ms for CIFAR10-NAS, 3 orders of magnitude lower than the MAX78000 (21.24 ms) and beating the HX-WE2 (0.12 ms). Combined with its short initialization latency (0.22 - 0.28 ms), the NXP-MCXN947, with further model inference optimization, could be well-suited to applications requiring frequent model switching or intermittency.
@@ 4.3. Task-Specific Considerations
The detailed memory I/O timing data provides additional insights into how different platforms handle model loading. The MAX78000’s long memory I/O times (8.84 - 26.53 ms) are more suitable for persistent model deployment. In contrast the HX-WE2’s low-latency memory I/O (0.03 - 1.11 ms) but longer initialization times (2.56 - 2.60 ms) are ideal for continuous inference or dynamic model switching.
Operational Modes and Power Profiles The ability to support different operational modes significantly impacts a platform’s suitability for specific applications. The MAX78000 displays high power variation between idle (10.87 - 13.21 mW) and inference (21.13 - 81.67 mW) states; hence, power gating mechanisms could be effectively leveraged in duty-cycled applications. Other platforms demonstrate high power consumption even when idle, which could limit their use in extremely power-constrained scenarios.
@@ 4.4. Summary of Results
The MAX78000 $\mu$NPU, with its Cortex-M4 CPU active, displays the best overall efficiency with NPU initialization included, delivering consistent sub-30ms end-to-end latency across all models. However, its performance is primarily memory-bound, spending up to 90% of execution time on memory I/O operations. The HX-WE2 platform achieves an average end-to-end $\sim$1.93x speedup over MAX78000 but with $\sim$3.13x higher power consumption. General-purpose MCUs demonstrate significantly lower efficiency, empirically validating the advantage of having dedicated neural hardware. Excluding initialization overhead ($i.e.$, for applications requiring continuous operation), the MILK-V ranks highest in terms of efficiency, with its large idle power draw outweighed by fast end-to-end inference latency.
@@ 4.6. Practical Recommendations
For Energy-Efficiency: The MAX78000 consistently outperforms other $\mu$NPU platforms in terms of energy-efficiency (when including NPU initialization), making it particularly well-suited for battery-powered applications. For extended battery life, consider leveraging its ability to dynamically power-gate portions of the system during idle periods.
For Latency-Critical Applications: The HX-WE2 platform offers low-latency with fast NPU initialization, memory I/O, and inference itself, making it best suited for applications requiring responsive model switching, real-time adaptation to changing conditions, or intermittent/duty-cycled operation. However, for latency-critical applications where power consumption isn’t a major constraint and workloads don’t require frequent NPU initialization/deinitialization, the MILK-V could also be suitable. Further work could explore the performance of other similar SoC platforms (luckfox, ; k230, ).
@@
For Security with Balanced Performance: Being Arm-based, the NXP-MCXN947 CPU includes Arm’s TrustZone (trustzone, ), enabling hardware isolation between secure and non-secure enclaves. Consequently, its quick initialization (0.22 – 0.28 ms), memory I/O (0.05 - 0.51 ms), and reasonable inference latency, may make it suitable for security-centric applications without extreme constraints in any one dimension. Future work could explore extending its secure execution environment to integrate $\mu$NPU acceleration via I/O passthrough, enabling protected yet efficient NN inference.

### v2

@@ Abstract.
In this paper, we present the first comparative evaluation of a number of commercially-available $\mu$NPUs, as well as the first independent benchmarks for several of these platforms. We develop and open-source a model compilation framework to enable consistent benchmarking of quantized models across diverse $\mu$NPU hardware. Our benchmark targets end-to-end performance and includes model inference latency, power consumption, and memory overhead, alongside other factors. The resulting analysis uncovers both expected performance trends as well as surprising disparities between hardware specifications and actual performance, including $\mu$NPUs exhibiting unexpected scaling behaviors with increasing model complexity. Our framework provides a foundation for further evaluation of $\mu$NPU platforms alongside valuable insights for both hardware designers and software developers in this rapidly evolving space.
@@ 3.1. Hardware
Figure 2. a visualization of the various $\mu$NPUs used in our benchmark, and how they compare in terms of GOPS, peak power draw observed, and theoretical efficiency (GOPS/mW).
@@ 3.2. Models
CIFAR10-NAS: the optimal CNN model generated by neural architecture search (NAS) for the CIFAR-10 dataset, combining diverse convolutional units, pooling layers, and unique connectivity patterns. The model was generated using the Once-for-All (OFA) NAS framework, a weight-sharing-based framework that decouples search and training by constructing a supernet model from which various hardware-specific subnet models can be derived (onceforall, ). This is our largest model, with 74.3 Million MACs (MMACs) and 36.4 Million FLOPs (MFLOPs). Trained on the CIFAR-10 dataset, with 3x32x32 input size and 10-class output.
@@ 3.2.1. Ensuring Model Uniformity
We encountered substantial variability in operator support across the benchmark platforms. The NXP-MCXN947’s eIQ Neutron NPU lacks native support for softmax operations, for example, necessitating its implementation as a CPU post-processing step for relevant models. Similarly, operations associated with non-maximum supression (NMS) in the YoloV1 model were inconsistently supported across platforms, requiring us to also move the entire NMS operation to CPU post-processing. This explains the unusual multi-component output shape of our YoloV1 model (see Table 2). The benchmark platforms also outline varying levels of support for operator compatibility. The MAX78000, for example, only supports 1D convolution with kernel sizes 1 to 9 and 2D convolution with kernel sizes of 1 by 1 or 3 by 3. Unsupported operations will fall back to CPU execution and incur latency penalties.
@@ 3.3. Evaluation Metrics
3.4. Performance Breakdown
@@
NAS 1.10±0.002 0.85±0.001 0.10±0.002 1.07±0.002 0.79±0.007 0.83±0.006 0.01±0.001 0.03±0.001 0.01±0.001
ResNet 1.24±0.003 0.85±0.002 0.17±0.002 1.97±0.003 0.85±0.006 0.84±0.019 0.01±0.001 0.06±0.001 0.02±0.001
SimpleNet 2.29±0.006 1.65±0.003 0.16±0.005 2.10±0.004 0.89±0.006 0.99±0.006 0.01±0.001 0.07±0.001 0.02±0.001
Autoenc 3.92±0.014 2.75±0.008 1.12±0.028 36.95±0.002 3.57±0.035 3.06±0.038 0.01±0.001 3.48±0.082 0.32±0.001
YoloV1_small 2.27±0.004 1.76±0.003 0.20± 0.005 1.83±0.006 0.73±0.009 0.81±0.008 0.01±0.001 0.05±0.001 0.01±0.001
@@
CIFAR10-NAS 1.11±0.002 0.85±0.001 0.11±0.002 1.09±0.002 0.98±0.009 1.04±0.008 2.75±0.281
ResNet 1.24±0.003 0.85±0.002 0.22±0.001 2.01±0.002 1.08±0.008 1.05±0.024 9.31±2.567
SimpleNet 2.30±0.006 1.66±0.003 0.21±0.007 2.13±0.004 1.13±0.008 1.29±0.010 4.13±0.459
Autoenc 3.94±0.014 2.78±0.008 6.25±0.203 47.06±1.956 22.45±0.392 12.97±0.232 13.74±1.953
YoloV1_small 2.27±0.004 1.76±0.003 0.23±0.005 1.86±0.007 0.91±0.013 1.03±0.011 5.75±0.770
@@ 4. RESULTS & DISCUSSION
Table Supplementary Material, which can found in the supplementary material, details our full latency and power measurements across each stage, model, and platform.
@@ 4.1. Power and Efficiency Breakdown
The MAX78000 (C) with Cortex-M4 CPU active demonstrates the best overall efficiency across evaluated models when NPU initialization overhead is considered, with consistent ¡30ms end-to-end latency. The MAX78000 (R) with RISC-V CPU lags slightly behind. This aligns with previous standalone benchmarks (moss2022ultra, ).
The NXP-MCXN947 also achieves consistent sub-30ms latency, with its fast initialization and memory I/O offsetting the impact of (moderately) slower inference latency, delivering comparable (and in some cases, improved) efficiency despite its lower-throughput accelerator.
@@
Notably, the power-hungry but low-latency HX-WE2 platform, with Arm Corstone-300 (Cortex-M55 & Ethos-U55 NPU), consistently beats the MAX78000 (C/R) in terms of end-to-end latency across the various models, due to the latter’s large memory I/O overhead. The HX-WE2 (S/P) demonstrates average $\sim$1.93x and $\sim$3.07x speedup in end-to-end latency over the MAX78000 (C) and (R) respectively, but $\sim$3.13x and $\sim$3.33x increase in average power consumption. We find the Vela Performance-optimized models, for the HX-WE2, generally achieve slightly lower latency than the Size-optimized models. However, their efficiency gain diminishes with model complexity – efficiency on Performance-optimized YoloV1 is lower than on its Size-optimized variant.
@@
The general-purpose MCUs without dedicated neural hardware – the STM32H7A3ZI and ESP32s3 – demonstrate significantly lower efficiency across all models. This result empirically validates the advantage neural hardware provides for performing on-device inference in constrained environments, with up to 2 orders of magnitude improvement in end-to-end latency in some cases. However, the STM32H7A3ZI’s power consumption during inference (54.91 - 56.11 mW) is comparable to or lower than MAX78000 (C/R) for some models. This is particularly evident for the Autoencoder model, where the STM32H7A3ZI achieves a surprisingly competitive 3.483 $I_{mJ}$ – comparable to the best-performing platforms in our suite. This anomaly is likely attributable to the STM32H7A3ZI’s relatively efficient Cortex-M7 core when operating on the Autoencoder’s simple computational structure (0.5455 MMACs). In contrast, the ESP32 consistently exhibits high inference power consumption (129.74 - 157.17 mW) and latency (7.11 – 536.22 ms), despite its advertised support for CPU-accelerated tensor operations. Altogether, while general-purpose MCUs can achieve reasonable efficiency for simple models, they quickly become impractical for more complex NNs.
@@
The MILK-V, our RISC-V SoC, also demonstrates low efficiency across all models, due to its NPU initialization overhead. We observe a different story, however, if initialization overhead is removed from consideration ($i.e.$, for continuous operation). Without initialization, the MILK-V ranks highest for efficiency across almost all benchmark models. Notably, despite a large idle power draw, it achieves blazingly fast inference times (0.17 - 0.61 ms). Thus, for applications where power consumption isn’t a major constraint and workloads don’t require frequent NPU initialization/deinitialization, but low-latency and a compact form factor are key, the MILK-V could prove effective.
@@
Fig. 7 details the power consumption breakdown across all evaluated platforms; among these, the MAX78000 (10.87–80.41 mW) and NXP-MCXN947 (22.91–36.69 mW) exhibit the lowest power draw across the benchmark models, with the NXP showing the lowest variance in peak power across the execution stages, enabling more reliable energy budgeting.
@@
Beyond peak power, idle power consumption is another key consideration for low-power deployments, particularly if workloads run infrequently – idle power also varies significantly across our benchmark platforms. The MAX78000 demonstrates the lowest idle power of the various $\mu$NPU platforms (10.87 mW with RISC-V and 13.21 mW under Cortex-M4). The HX-WE2 platform ranks highest (89.09 mW), raising concerns about its applicability in extremely power-constrained scenarios (such as ones in which long idle durations dominate overall energy usage).
@@ 4.2.2. Memory I/O
The significant memory I/O latency across all models on the MAX78000 forms an obvious inference bottleneck, with an average of 6.10x and 9.80x (Cortex-M4 and RISC-V) longer spent on memory I/O than actual inference ($e.g.$, 44.89 ms vs. 2.96 ms for ResidualNet with the MAX78k (R), meaning over 90% of end-to-end inference time is dedicated to memory operations rather than computation). This implies the MAX78000’s performance is largely memory-bound, and aligns with previous standalone benchmarks (moss2022ultra, ). Notably, memory I/O operations are more efficient on the MAX78000’s Cortex-M4 CPU than its RISC-V one. In contrast, memory I/O operations introduce negligible overhead across the other benchmark platforms with shared SRAM.
@@ 4.3. Task-Specific Considerations
The detailed memory I/O timing data provides additional insights into how different platforms handle model loading. The MAX78000’s long memory I/O times (8.84 - 26.53 ms) are more suitable for persistent model deployment. In contrast the HX-WE2’s comparatively large flash memory and low-latency memory I/O (0.03 - 1.11 ms), but longer initialization times (2.56 - 2.60 ms), are ideal for continuous inference or dynamic model switching.
Operational Modes and Power Profiles The ability to support different operational modes significantly impacts a platform’s suitability for specific applications. The MAX78000 displays high power variation between idle (10.87 - 13.21 mW) and inference (21.13 - 81.67 mW) states; hence, power gating mechanisms could be effectively leveraged in duty-cycled applications.
@@ 4.4. Summary of Results
The MAX78000 $\mu$NPU, with its Cortex-M4 CPU active, offers the best overall efficiency, with NPU initialization considered, delivering consistent sub-30ms end-to-end latency across all models. However, its performance is primarily memory-bound, spending up to 90% of execution time on memory I/O operations. The HX-WE2 platform achieves an average end-to-end $\sim$1.93x speedup over MAX78000 but with $\sim$3.13x higher power consumption. The NXP-MCXN947 also offers relatively comparable (¡30ms) end-to-end latency, with fast initialization and memory I/O; despite its lower computational throughput, it exhibits high efficiency on our lower-complexity, memory-light benchmark models.
General-purpose MCUs demonstrate significantly lower efficiency, empirically validating the advantage of having dedicated neural hardware.
Excluding initialization overhead ($i.e.$, for applications requiring continuous operation), the MILK-V ranks overall highest in terms efficiency, with its large idle power draw outweighed by fast end-to-end inference latency. Notably, the NXP-MCXN947 ranks highest, both with and without NPU initialization, for the Autoencoder model, delivering over 2× efficiency gains on its nearest competitor.
@@ 4.6. Practical Recommendations
For Energy-Efficiency: The MAX78000 largely outperforms other $\mu$NPU platforms in terms of energy-efficiency (when including NPU initialization), making it particularly well-suited for battery-powered applications. For extended battery life, consider leveraging its ability to dynamically power-gate portions of the system during idle periods.
For Latency-Critical Applications: The HX-WE2 platform offers low-latency with fast NPU initialization, memory I/O, and inference itself, making it best suited for applications requiring responsive model switching, real-time adaptation to changing conditions, or intermittent/duty-cycled operation. The NXP-MCXN947 also achieves low end-to-end inference latency at a significantly lower power budget, making it ideal for power-constrained workloads. Meanwhile, for latency-critical yet space-constrained applications, where power consumption isn’t a major constraint and workloads don’t require frequent NPU initialization/deinitialization, more powerful SoC architectures, like the MILK-V, could also be suitable. Further work could explore the performance of other SoC platforms (luckfox, ; k230, ).
@@
For Security with Efficiency: Being Arm-based, the NXP-MCXN947 CPU includes Arm’s TrustZone (trustzone, ), enabling hardware isolation between secure and non-secure enclaves. With its competitive end-to-end inference latency (0.96-25.95 ms) and low power draw (22.91–36.69 mW), it may be suitable for security-centric applications without extreme constraints in any one dimension. Future work could explore extending its secure execution environment to integrate $\mu$NPU acceleration via I/O passthrough, enabling protected yet efficient NN inference.